for teminals A and B the circuit is equivalent to:


Grade:12th Pass

2 Answers

Aman Bansal
13 years ago

Dear Aditya,

The correct answer is B

BEST OF LUCK..!!!!

Plz approve the answer....!!

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL

Sanjeev Malik IIT BHU
13 years ago

consider the currnt through the closed loop

i= 10/(100+100)=1/20

so the potential across 100 ohm resistor between the joints (or u can say A and B as free ends has no current)

= 1/20*100=5 V

so that passage equivalent to 5V

so correct ans is A.

 

 

please approve......